solve the following system of equations by the substitution method:
4x=3y+30 3y-5x=-33
after substituting, find the remaining equation and the y coordinate of the solution.

Answer by checkley71(8403)

3Y-5X=-33 OR 3Y=5X-33 NOW REPLACE 3Y IN THE OTHER EQUATION BY (5X-33) & SOLVE FOR X:
4X=(5X-33)+30
4X=5X-33+30
4X-5X=-3
-X=-3
X=-3/-1
X=3
3Y=5*3-33
3Y=15-33
3Y=-18
Y=-18/3
Y=-6 ANSWER.